# MTLoc: A Confidence-Based Source-Free Domain Adaptation Approach For Indoor Localization

### v1

@@ Abstract
Various deep learning models have been developed for indoor localization based on radio-frequency identification (RFID) tags. However, they often require adaptation to ensure accurate tracking in new target operational domains. To address this challenge, unsupervised domain adaptation (UDA) methods have been proposed to align pre-trained models with data from target environments. However, they rely on large annotated datasets from the initial domain (source). Source data access is limited by privacy, storage, computational, and transfer constraints. Although many source-free domain adaptation (SFDA) methods address these constraints in classification, but applying them to regression models for localization remains challenging. Indeed, target datasets for indoor localization are typically small (few features and samples) and noisy, and adapting regression models requires high-confidence target pseudo-annotation to avoid over-training. In this paper, a specialized mean-teacher method called MTLoc is proposed for SFDA. MTLoc updates the student network using noisy data and teacher-generated pseudo-labels. The teacher network maintains stability through exponential moving averages. To further ensure robustness, the teacher’s pseudo-labels are refined using $k$-nearest neighbor correction. MTLoc allows for self-supervised learning on target data, facilitating effective adaptation to dynamic and noisy indoor environments. Validated using real-world data from our experimental setup with INLAN Inc., our results [^1] Code and datasets: https://github.com/negarmehregan/Indoor-Localization/tree/main show that MTLoc achieves high localization accuracy under challenging conditions, significantly reducing localization error compared to baselines, including the state-of-the-art adversarial UDA approach with access to source data.
Index Terms:
Deep learning, unsupervised domain adaptation, source-free domain adaptation, mean-teacher models, indoor localization.
I Introduction
@@
Figure 1: Illustration of an indoor localization system for tracking objects and individuals in real-time having three receivers and a transmitter by employing tags such as UWB, RFID, BLE.
@@
As GPS signals are proven to be ineffective for localization due to signal deflection and multipath-related issues [6], different technologies are introduced, with qualifications varying based on the type of tag applied. Among others, ultra-wideband (UWB) tags [7, 8] with high-precision accuracy and low power usage, radio frequency identification (RFID) tags [9, 10] with lower cost and long-range communication, and Bluetooth low energy (BLE) tags [11, 12], a medium range, energy-efficient technology, are utilized in localization models. Tag-based tracking models are employed to monitor key assets’ location, ensure safety, and localize objects for healthcare applications[13, 14, 15]. Fig. 1 illustrates the application of an indoor localization system, in which a signal is passed from the transmitter to the receiver, after reaching the tagged object. A typical indoor positioning system incorporates receivers, transmitters, and tags for data collection. In various studies, through using receivers, different types of data were acquired from tags including received signal strength [16, 17, 18], Phase, Channel Impulse Response [19, 20], channel state information (CSI) [21, 22], angle of arrival [23, 24, 25], and radio frequency [26, 27]. However, the inaccuracies caused by signal attenuation, and distortion, remain a challenge, particularly in complex environments in which multipath components and non-line-of-sight effects stand out as the most concerning[28, 20]. To address these issues, various deep learning-based techniques are proposed for indoor localization.
Several AI-based techniques are being developed to address indoor localization problems. For instance, with MFFALoc [21] CSI fingerprints are fed into a convolutional neural network (CNN), tackling multipath and non-line-of-sight effects. Furthermore, Rao et al. [21] remove environmental noise by utilizing waveform diversity and detection, and implement a feature fusion system to merge the multi-antenna feature matrices, resulting in fine-grained fingerprints. Some studies, such as Chen et al. [29], apply 1D-CNN for RF-based environment classification, while others, like Yan et al. [30], use 2D deep CNNs for CSI feature extraction to improve localization accuracy. Se-Loc [31], is another CNN-driven approach with a security-enhanced semi-supervised indoor localization technique by integrating a denoising autoencoder (AE), aiding in improving fingerprinting localization. This model has a pearson correlation coefficient-based access point selection system to eliminate the contaminated access points. Certain research initiatives, such as in [32], use AE for spatial feature extraction. Likewise, in [33] and [28], a variational AE is being applied to extract key position-related features and reduce environmental noise to solve localization and multipath challenges. Some researchers merge recurrent neural networks with AE such as  [34], where a stacked AE, extracting key features, is combined with an attention-based long short-term memory network, enhancing its localization accuracy. On the other hand, Liu et al.[35] uses gated recurrent units for feature extraction, while Chen et al.[36] apply a deep LSTM with a local feature extractor to process RSSI fingerprints, capturing temporal dependencies through a regression layer for localization. Some use generative models for data augmentation [37, 38, 39], improving localization accuracy.
@@
Previous studies discover the impact of environmental change on the data [40, 41]. As mentioned by [42], CSI data is affected by any changes in the environment, hence, a localization system might not perform well in a new environment. Therefore, they propose an unsupervised domain adaptation (UDA) method in which the source domain’s model is fine-tuned to be applicable in a new distinct but associated domain, where the ground truth data is not accessible. Given a pre-trained model utilizing labeled source data collected from Warehouse A, the objective is to evaluate its performance in a different environment, Warehouse B. The primary challenge lies in the fact that the new domain (Warehouse B) may produce substantially different object-related signals, possibly due to variations in the positions of transmitters and receivers. However, retraining the model is infeasible due to the lack of annotated data from Warehouse B. This scenario falls under the framework for UDA.
Currently, UDA approaches can be classified into three main categories [43]: a) domain distribution alignment [44, 45, 46] b) adversarial learning to extract domain-invariant representation [47, 48, 49], and c) model fine-tuning through self-supervision [50, 51, 52]. In the context of indoor localization, most UDA approaches rely on adversarial learning through the use of the gradient reversal layer (GRL), as Zhang et al. [53] utilizing a semi-supervised graph convolutional network based model with received signal strength indicator data, Xu et al. [54] with a gaussian moise infused co-teaching mechanism, Zhang et al. [55] in which dynamic adversarial adaptation networks is applied to eliminate any inconsistency in both global and local discriminators, and Prasad et al. [56], applying an AE for dimension reduction purposes as well as gradient reversal layer to create a domain invariant features. Additionally, discrepancy-based models, aligning source and target domain distributions are created, as shown in [57], in which the grassmann manifold is employed for dimensional reduction purposes while preserving the integrity of original data, and maximum mean discrepancy to ensure localization accuracy. Certain studies conducted hybrid mechanisms by using both generative and GRL-based approaches in their models, for instance, [58] and [42] in which generative models are used for data augmentation, using variational AE and adversarial AE respectively. In [58], the researcher applied a domain-adaptive classifier incorporating a feature extractor as well as a classification-reconstruction layer. Meanwhile, in [42], a CNN model serves as a feature extractor, with a gradient reversal layer-based discriminator and location predictor networks.
Nevertheless, as there are concerns regarding the data-privacy issues in UDA, especially, in case of having sensitive information in the source domain, as mentioned by [59], therefore, to avoid data leakage, source-free based models could be applied for localization purposes, aiding in achieving high accuracy across different domains while respecting the privacy concerns. On top of that, source-free (unsupervised) domain adaptation (SFDA) systems reduce computable costs and facilitate data storage, and transfer [60, 61]. Moreover, in certain cases, regulations prevent any data transfer, while some companies also aim to maintain the ownership of their data, thus, making source-free models more practical [61]. Furthermore, the scalability of SFDA models enables organizations to adapt accordingly to different target environments [61, 62]. SFDA is considered a way to create an efficient domain-invariant system as discussed by Fang et al. [60] who conducted research in this area, providing a detailed survey regarding image-based studies.
@@
To the best of our knowledge, no method for SFDA has been proposed for indoor localization. Of the current SFDA methods, the mean teacher framework is particularly well-suited for regression-based indoor localization. In practice, target datasets for indoor localization are often limited and noisy, making domain adaptation reliant on highly confident pseudo-labels. Researchers such as Yeonguk et al. [63] have used this technique due to its EMA-based stabilization, which produces stable pseudo labels through consistency regularization. Although it is a simple approach, it ensures efficient performance in source-free scenarios while leveraging self-supervision. Motivated by these characteristics, we introduce mean teacher localization (MTLoc) for SFDA, a self-supervised method leveraging a mean teacher framework. This approach uses a teacher-student architecture, where both networks are initialized with a localizer, pre-trained on source data. The student model is fine-tuned on noisy target domain data using pseudo labels generated by the teacher network, enabling effective adaptation without access to source data. To further improve performance, we propose an enhanced model, MTLoc with confidence, where the teacher’s uncertain pseudo-labels are corrected using $k$-nearest neighbors ($k$-NN), and the student model is trained exclusively on the most confident ones generated by the teacher network. This refinement improves the reliability of the adaptation by focusing on high-certainty predictions, reducing variations, thereby enhancing the robustness of the SFDA process.
@@
  • We conducted a comprehensive data collection process to develop three datasets—Ceiling (source), Cross, and Square (targets)—at the INLAN Inc. testing facility. This setup involved four receivers, a transmitter equipped with dipole antennas, and an RFID tag, providing a robust evaluation environment for the proposed method. MTLoc is validated on this real-world data and significantly outperforms the state-of-the-art adversarial UDA model.
@@
This paper is organized as follows. Section II provides a problem definition for SFDA, followed up by details on our domain adaptation baseline method. In section III, we provide specifics on our proposed confidence-based SFDA method. Section IV presents our methodology for experimental validation (i.e., data collection, protocol, performance measures), as well as our analysis of experimental results.
Notation and Conventions: We denote random variables using uppercase letters (e.g., $X,Y$) and their specific realizations using lowercase letters (e.g., $x,y$). The notation $p(x)$ refers to the probability distribution of the random variable $X$, while $p(x|y)$ represents the conditional distribution of $X$ given $Y$. The expectation of a function with respect to the joint distribution of all random variables is denoted by $\mathbb{E}[\cdot]$.
II Problem Definition and Baseline Overview
II-A Problem Formulation:
Consider an indoor localization framework consisting of $N$ receivers and a single transmitter. The transmitter emits signals that interact with objects within the environment, and these signals are subsequently captured by the receivers. The objective is to develop an object localization system by analyzing the signals recorded by the receivers. Let $Z$ denote the random variable representing the signals recorded by the receivers, and let $(X,Y)$ represent the joint random variables corresponding to the true location associated with the recorded signal $Z$. The goal is to learn an object localizer defined as the mapping $\mathcal{M}(\cdot;\theta):\mathcal{Z}\rightarrow\mathcal{X}\times\mathcal{Y}$ parameterized by $\theta$, which produces an approximation $(\hat{X},\hat{Y})$ of the true location for each observed signal $Z$, i.e $(\hat{X},\hat{Y})=\mathcal{(}Z;\theta)$. To optimize the parameters $\theta$ of the localizer, we employ Bayesian optimization by maximizing the log-likelihood of the true location given the estimated location as follows:
$\theta^{*}=\arg\max_{\theta}\log p_{X,Y\mid\hat{X},\hat{Y}}(x,y\mid\hat{x},\hat{y}),$ (1)
@@ II-A Problem Formulation:
Gaussian Error Model[64]: Assuming the localization errors $(X-\hat{X},Y-\hat{Y})$ follow a bivariate Gaussian distribution with zero mean and covariance matrix $\Sigma$, the negative log-likelihood function becomes:
$-\log p(x,y\mid\hat{x},\hat{y})\propto\begin{pmatrix}x-\hat{x}\\ y-\hat{y}\end{pmatrix}^{\top}\Sigma^{-1}\begin{pmatrix}x-\hat{x}\\ y-\hat{y}\end{pmatrix}$ (2)
@@
Laplacian Error Model[64]: Alternatively, considering the localization errors follow a bivariate Laplacian distribution and assuming independence between error in $X$ and $Y$, the joint probability density function can be written as follows:
$p(x,y\mid\hat{x},\hat{y})=\left(\frac{1}{\sqrt{2}b}\right)^{2}\exp\left(-\frac{|x-\hat{x}|+|y-\hat{y}|}{\sqrt{2}b}\right)$ (4)
@@
$-\log p(x,y\mid\hat{x},\hat{y})=\frac{1}{\sqrt{2}b}\left(|x-\hat{x}|+|y-\hat{y}|\right)+2\log(\sqrt{2}b)$ (5)
@@
II-B Baselines:
In this study, we considered several baselines including a source-only pre-adaptation model and an oracle approach trained on labeled target datasets [65, 66]. Among UDA approaches created for indoor localization, adversarial UDA and discrepancy-based models are prominent, with the state-of-the-art focusing predominately on the use of GRL-based adversarial UDA techniques. Thereby, we evaluated an adversarial UDA method that leverages a GRL-based model. This approach is inspired by FreeLoc [42], an adaptation of Domain-Adversarial Neural Network (DANN) [67], which facilitates domain alignment through adversarial training, and serves as another baseline. As illustrated in Fig. 2, the framework assumes access to a labeled source dataset, $\mathbf{D}_{s}=\{z_{s}^{(i)},(x_{s}^{(i)},y_{s}^{(i)})\}_{i=1}^{N_{s}}$ with distribution $p^{s}_{X,Y,Z}$, and an unlabeled target dataset, $\mathbf{D}_{t}=\{z_{t}^{(j)}\}_{j=1}^{N_{t}}$ with distribution $p^{t}_{Z}$. Here, $z$ represents the signals received by the receivers, while $(x,y)$ denotes the object location in the corresponding domain.

### v2

@@ Abstract
Indoor localization using radio-frequency identification (RFID) has benefited from deep learning, yet models trained in a labeled source environment often degrade when deployed in a different, unlabeled target environment. Unsupervised domain adaptation (UDA) aims to mitigate this distribution shift by aligning a source-trained model with target-domain data. In practice, the source dataset is frequently unavailable at adaptation time due to privacy and resource constraints. This motivates source-free domain adaptation (SFDA); however, most SFDA methods have been developed for classification, and extending them to indoor localization (regression) is challenging, especially when target datasets are small and noisy. Motivated by above limitations, we introduce MTLoc, a source-free mean-teacher approach for indoor localization. MTLoc includes a student and a teacher network: the student network is updated using noisy target data with teacher-generated pseudo-labels. The teacher network maintains stability through exponential moving averages. To further ensure robustness, we propose a correction mechanism in which the teacher’s pseudo-labels are refined using $k$-nearest neighbor correction. MTLoc allows for self-supervised learning on target data, facilitating effective adaptation to dynamic and noisy indoor environments. Validated using real-world data from our experimental setup with INLAN Inc., our results1 show that MTLoc achieves high localization accuracy under challenging conditions, significantly reducing distance error compared to baselines. On average, it reduces MAE by $20.0\%$ on Cross and $22.5\%$ on Square datasets. With confidence correction, these improvements reach $23.9\%$ and $28.2\%$ respectively.
{IEEEkeywords}
Source-free domain adaptation, Indoor localization, Regression, Deep learning, Pseudo-labeling.
11footnotetext: Code and datasets: https://github.com/negarmehregan/Indoor-Localization/tree/main
[…]
NOMENCLATURE
$X,Y$ Random variables
$x,y$ Realizations for $X,Y$
$p(x)$ Probability distribution of $X$
$p(x\mid y)$ Conditional distribution of $X$ given $Y$
$\mathbb{E}[X]$ Expectation of the random variable $X$
EMA ($\alpha$) Exponential moving average with parameter $\alpha$
GRL Gradient Reversal Layer
$k$-NN $k$-nearest-neighbor
1 Introduction
[…]
@@ 1 Introduction
As GPS signals are proven to be ineffective for localization due to signal deflection and multipath-related issues [10706359], different technologies are introduced, with qualifications varying based on the type of tag applied. Notably, ultra-wideband (UWB) tags [10756560, huang2023indoor] with high-precision accuracy and low power usage, radio frequency identification (RFID) tags [7842591, yang2014tagoram] with lower cost and long-range communication, and Bluetooth low energy (BLE) tags [9075151, 9343321], a medium range, energy-efficient technology, are used in localization models.
@@
Tag-based tracking models are used to monitor key assets’ location, ensure safety, and localize objects for healthcare applications [6587041, 5162266, 5304900]. A typical indoor positioning system includes receivers, transmitters, and tags for data collection, where a signal is transmitted to the receiver after interacting with the tagged object. In various studies, through using receivers, different types of data were acquired from tags including received signal strength [10000400, 10190065, 10705937], Phase, Channel Impulse Response [9903809, 10006709], channel state information (CSI) [Rao2024, 10838284], angle of arrival [10472153, 10260273, 9676583], and radio frequency [9244574, 9555623]. However, the inaccuracies caused by signal attenuation and distortion remain a challenge, particularly in complex environments in which multipath components and non-line-of-sight effects stand out as the most concerning [Wang2024, 10006709]. To address these issues, various deep learning-based techniques are proposed for indoor localization.
@@
Previous studies highlight the impact of environmental change on data [chen2020FiDO, li2021DAFI]. As mentioned by [Yan2024], CSI data is affected by any changes in the environment, hence, a localization system might not perform well in a new environment. Therefore, unsupervised domain adaptation (UDA) methods have been proposed, where the source domain model is fine-tuned to be applicable in a distinct but related target domain without access to ground truth. Given a pre-trained model utilizing labeled source data collected from Warehouse A, the objective is to evaluate its performance in a different environment, Warehouse B. The primary challenge lies in the fact that the new domain (Warehouse B) may produce substantially different object-related signals, possibly due to variations in the positions of transmitters and receivers. Retraining the model is infeasible due to the lack of annotated data from Warehouse B, falling under the framework for UDA.
Nevertheless, UDA raises concerns regarding data privacy, particularly when sensitive information is present in the source domain [9200758]. To avoid data leakage, source-free domain adaptation (SFDA) has been proposed, which enables robust adaptation without access to source data. SFDA not only respects privacy concerns but also reduces computational costs and facilitates storage and transfer [Fang2023, 10452835]. Additionally, regulatory constraints or corporate policies often restrict data transfer, further motivating SFDA as a practical approach [10452835]. The scalability of SFDA also supports adaptation across diverse environments [10452835, 10078842]. While most SFDA methods target classification, TASFAR [10597720] is a recent regression-based approach that estimates the target label distribution by modeling predictive uncertainty via a label-density map. It generates pseudo-labels for target samples to finetune the source model.
@@
To the best of our knowledge, no method for SFDA has been proposed for indoor localization. Of the current SFDA methods, the mean teacher framework is particularly well-suited for regression-based indoor localization. In practice, target datasets for indoor localization are often limited and noisy, making domain adaptation reliant on highly confident pseudo-labels. Researchers such as Yeonguk et al. [Yu2024] have used this technique due to its EMA-based stabilization, which produces stable pseudo labels through consistency regularization. Although it is a simple approach, it ensures efficient performance in source-free scenarios while leveraging self-supervision.
Motivated by these characteristics, we introduce mean teacher localization (MTLoc) for SFDA, a self-supervised method leveraging a mean teacher framework. This approach uses a teacher-student architecture, where both networks are initialized with a localizer pre-trained on source data. The student model is fine-tuned on noisy target domain data using pseudo labels generated by the teacher network, enabling effective adaptation without access to source data. To further improve performance, we propose an enhanced model, MTLoc with confidence, where the teacher’s uncertain pseudo-labels are corrected using $k$-nearest neighbors ($k$-NN), and the student model is trained exclusively on the most confident ones. This refinement improves reliability by focusing on high-certainty predictions, reducing variations, and enhancing robustness.
@@
  • A comprehensive data collection process was conducted to develop three datasets—Ceiling (source), Cross, and Square (targets)—at the INLAN Inc. testing facility. This setup involved four receivers, a transmitter equipped with dipole antennas, and an RFID tag, providing a robust evaluation environment for the proposed method. MTLoc is validated on this real-world data and significantly outperforms the state-of-the-art UDA and SFDA models.
@@
This paper is organized as follows. Sections 2 provides related work. Section 3 defines the SFDA problem and baseline. Section 4 presents our proposed confidence-based SFDA method. Section 5 details data collection, experimental protocol, performance measures, and results analysis.
2 Related Work
Over the years, several approaches have been proposed for indoor localization. Some rely on kernel-based methods, such as Zhang et al. [Zhang2021IMKELM], who developed an integrated multiple-kernel extreme learning machine that uses temporal and spatial CSI features for device-free localization in cluttered environments. However, AI-based techniques have become the dominant trend. For instance, with MFFALoc [Rao2024], CSI fingerprints are fed into a convolutional neural network (CNN), tackling multipath and non-line-of-sight effects. Rao et al. [Rao2024] further remove environmental noise by utilizing waveform diversity and detection, and implement a feature fusion system to merge multi-antenna feature matrices, resulting in fine-grained fingerprints. For example, MFFALoc [Rao2024] use a CNN on CSI fingerprints to tackle multipath and non-line-of-sight effects. In addition, Rao et al. [Rao2024] improve fingerprint quality by using waveform diversity and detection to suppress environmental noise, and by fusing multi-antenna feature matrices to construct fine-grained representations. In parallel, Zhang et al. [10930823] address domain adaptation by using online learning, where fingerprint augmentation and a sequential extreme learning machine with a forgetting mechanism allow the model to adapt dynamically to environmental changes. Some studies, such as Chen et al. [9082193], apply 1D-CNN for RF-based environment classification, while others, like Yan et al. [9541367], use 2D deep CNNs for CSI feature extraction to improve localization accuracy. Se-Loc [Ye2023] is another CNN-driven approach with a security-enhanced semi-supervised indoor localization technique, integrating a denoising autoencoder (AE) to improve fingerprinting. This model employs a Pearson correlation coefficient-based access point selection system to eliminate contaminated access points. Cao et al. [113645] proposed a confidence-aware mean teacher framework for metallographic image semantic segmentation, which adaptively calibrates pseudo-labels with class-aware thresholds and improves data augmentation through confidence-driven patch mixing.
Some research works, such as [Ruan2023], use AE for spatial feature extraction. Likewise, in [Li2023] and [Wang2024], a variational AE is applied to extract key position-related features and reduce environmental noise to solve localization and multipath challenges. Some works merge recurrent neural networks with AE, such as  [Ayinla2024], where a stacked AE is combined with an attention-based long short-term memory network, enhancing localization accuracy. Liu et al.[Liu2024] use gated recurrent units for feature extraction, while Chen et al.[8733822] apply a deep LSTM with a local feature extractor to process RSSI fingerprints, capturing temporal dependencies through a regression layer. Generative models have also been explored for data augmentation [Junoh2024, Li2021, Wu2023], improving localization accuracy.
In fault diagnosis settings, Wang et al. [wang2025domain] provide a comprehensive survey of domain adaptation, focusing on both homogeneous methods, including metric learning, adversarial learning, and feature reconstruction, along with heterogeneous settings, source-free, domain generalization, partial, open-set, and universal domain adaptation methods.
Currently, UDA approaches can be classified into three main categories [Li2024]: (a) domain distribution alignment [Long2015, Cui2020, Shui2023], (b) adversarial learning to extract domain-invariant representation [Long2018, Tzeng2017, Li2022], and (c) model fine-tuning through self-supervision [Lian2019, Zou2018, Li2022Divergence]. In indoor localization, most UDA approaches rely on adversarial learning through gradient reversal layers (GRL). Examples include Zhang et al. [Zhang2023], who use a semi-supervised graph convolutional network with RSSI data; Xu et al. [Xu2024], who propose a Gaussian noise–infused co-teaching mechanism; Zhang et al. [10745164], who develop dynamic adversarial adaptation networks to eliminate inconsistencies in global and local discriminators; and Prasad et al. [Prasad2024], who apply an AE for dimension reduction alongside a GRL to create domain-invariant features.
Discrepancy-based models aligning source and target distributions have also been explored, such as [Zhang2024a], which uses the Grassmann manifold for dimensionality reduction while preserving data integrity, combined with maximum mean discrepancy for accuracy. Hybrid mechanisms combining generative and GRL-based approaches also exist, such as [Chen2022] and [Yan2024], where generative models (variational AE and adversarial AE, respectively) are used for augmentation. In [Chen2022], a domain-adaptive classifier includes a feature extractor with classification-reconstruction layers, while in [Yan2024], a CNN feature extractor is coupled with GRL-based discriminator and predictor networks.
As discussed earlier, SFDA overcomes the limitations of UDA regarding privacy, storage, and transfer constraints. Fang et al. [Fang2023] provide a detailed survey of SFDA in image-based studies, highlighting its efficiency and scalability. However, no SFDA methods have yet been applied to indoor localization, leaving an open gap that this paper addresses.
3 Problem Definition and Baseline Overview
3.1 Problem Formulation:
[…]
@@ 3.1 Problem Formulation:
Gaussian Error Model[bishop2006pattern]: Assuming the localization errors (XX^,YY^)(X-$\hat${X},Y-$\hat${Y}) follow a bivariate Gaussian distribution with zero mean and covariance matrix $\Sigma$, the negative log-likelihood function becomes:
$\log p(x,y\mid\hat{x},\hat{y})\propto\begin{pmatrix}x-\hat{x}\\ y-\hat{y}\end{pmatrix}^{\top}\Sigma^{-1}\begin{pmatrix}x-\hat{x}\\ y-\hat{y}\end{pmatrix}$. (2)
@@
Laplacian Error Model[bishop2006pattern]: Alternatively, considering the localization errors follow a bivariate Laplacian distribution and assuming independence between error in $X$ and $Y$, the joint probability density function can be written as follows:
$p(x,y\mid\hat{x},\hat{y})=\left(\frac{1}{\sqrt{2}b}\right)^{2}\exp\left(-\frac{|x-\hat{x}|+|y-\hat{y}|}{\sqrt{2}b}\right),$ (4)
@@
$\log p(x,y\mid\hat{x},\hat{y})=\frac{1}{\sqrt{2}b}\left(|x-\hat{x}|+|y-\hat{y}|\right)+2\log(\sqrt{2}b).$ (5)
@@
3.2 Baselines:
In this study, we considered several baselines including a source-only pre-adaptation model and an oracle approach trained on labeled target datasets [varailhon2024sourcefreedomainadaptationyolo, Belal_2024_WACV]. Among UDA approaches created for indoor localization, adversarial UDA and discrepancy-based models are prominent, with the state-of-the-art focusing predominately on the use of GRL-based adversarial UDA techniques. Thereby, we evaluated an adversarial UDA method that leverages a GRL-based model. This approach is inspired by FreeLoc [Yan2024], an adaptation of Domain-Adversarial Neural Network (DANN) [Ganin2015DANN], which facilitates domain alignment through adversarial training, and serves as another baseline. As illustrated in Fig. 1, the framework assumes access to a labeled source dataset, $\mathbf{D}_{s}=\{z_{s}^{(i)},(x_{s}^{(i)},y_{s}^{(i)})\}_{i=1}^{N_{s}}$ with distribution $p^{s}_{X,Y,Z}$, and an unlabeled target dataset, $\mathbf{D}_{t}=\{z_{t}^{(j)}\}_{j=1}^{N_{t}}$ with distribution $p^{t}_{Z}$. Here, $z$ represents the signals received by the receivers, while $(x,y)$ denotes the object location in the corresponding domain.